• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name E. Schulte

    Log entry time 16:01:18 on November 03, 2008

    Entry number 246470

    keyword=Day Shift Summary, November 3, 2008

    This is a reconstruction of the shift log from the day shift, since I
    accidentally blitzed the other one...

    08:00 Shift starts debugging a high dead time problem with the DAQ. It
    turns out to be a problem with the BigBite wire chambers.

    08:30 The HRS-L Q3 had another cryo-related interlock fault. The liquid
    helium level recovered on its own and the magnet was successfully reset.

    09:40 A controlled access was requested to check on the BigBite wire
    chambers. It was decided that the thresholds would be increased to make
    the symptom of the dead time problems go away.

    10:30 Returned from access. Requested return to Beam Permit and beam
    delivery. We checked to beam positions with the BeO target and performed
    a harp scan.

    10:50 Begin Setup for the 3He running.

    11:13 The harp scan is performed after a delay due to beam problems.

    11:20 Begin first NMR measurement of the 3He target. This included
    setting up the target for longitudinal polarization. Polarization is
    about 47%.

    12:00 Requested beam delivery on 3He target from MCC. Requested 5uA with
    a 4x4 raster.

    13:45 Gas shed air flow and low pressure alarms. These turned out to be
    caused by a bad fan. J. Segal is looking into it.

    13:49 Begin second NMR measurement. Polarization is about 42%.

    13:57 Request return of beam delivery.

    15:40 Begin third NMR measurement. Polarization is aobut 37%.

    15:58 Request return of beam delivery. Start a new run 2753 in "setting
    6" mode.